# VortSDF: 3D Modeling with Centroidal Voronoi Tesselation on Signed Distance Field

### v1

@@ Abstract
Volumetric shape representations have become ubiquitous in multi-view reconstruction tasks. They often build on regular voxel grids as discrete representations of 3D shape functions, such as SDF or radiance fields, either as the full shape model or as sampled instantiations of continuous representations, as with neural networks. Despite their proven efficiency, voxel representations come with the precision versus complexity trade-off. This inherent limitation can significantly impact performance when moving away from simple and uncluttered scenes. In this paper we investigate an alternative discretization strategy with the Centroidal Voronoi Tesselation (CVT). CVTs allow to better partition the observation space with respect to shape occupancy and to focus the discretization around shape surfaces. To leverage this discretization strategy for multi-view reconstruction, we introduce a volumetric optimization framework that combines explicit SDF fields with a shallow color network, in order to estimate 3D shape properties over tetrahedral grids. Experimental results with Chamfer statistics validate this approach with unprecedented reconstruction quality on various scenarios such as objects, open scenes or human.
@@ 3.2 Coarse-to-fine Centroidal Voronoi Tesselation
Figure 3: The site locations are optimized using an approximated CVT algorithm that does not explicitly identifies the Voronoi cells (left) but consider the neighboring bisectors instead (right).
@@ 3.2.1 SDF-aware implicit CVT optimization
where $b(\mathbf{s_{i}},\mathbf{s_{j}},\mathbf{r})$ the distance from $\mathbf{s_{i}}$ to its bisector with $\mathbf{s_{j}}$ in the direction $\mathbf{r}$.
@@ 3.4.1 Normal Smoothing
$L_{reg}=0.5\sum_{t}(1-\cos(\mathbf{\nabla sdf_{t}},\mathbf{\nabla sdf_{t}}^{smooth}))^{2}.$
@@
where $w_{ek}$, $w_{reg}$ and $w_{tv}$ are weight factors.
@@ 4.1 Metrics
We evaluate the quality of the estimated geometry using the available ground truth 3D meshes and a point to mesh Chamfer distance. We compute these errors from the ground truth mesh to the predicted meshes to obtain accuracy $Acc$ and from the predicted mesh to the ground truth mesh to obtain completeness $Compl$. We clip the errors to a maximum distance of $0.1$ meter. Note that smaller values are better for these metrics.
@@ 4.2 Experiments on BlendedMVS dataset
We evaluate our method at the last $2$ levels of densification. Our method usually terminates with about $2$M points at lvl $5$ and about $500$K points at lvl $4$. In comparison VOXURF uses a uniform grid of $256\times 256\times 256$ voxels, which corresponds to about $16$M points.
@@
Dog $1.54$ $19.1$ $1.79$ $\mathbf{4.81}$ $\underline{0.98}$ $11.6$ 1.83/1.83/1.83 /$\mathbf{0.96}$ 18.72/18.72/18.72 /$\underline{9.88}$
Bear $6.10$ $\mathbf{11.25}$ $\underline{2.07}$ $70.8$ $3.17$ $44.08$ 2.20/2.20/2.20 /$\mathbf{1.41}$ $3/3$$\underline{26.3}$
Clock $1.34$ $11.8$ $1.11$ $3.29$ $\underline{0.95}$ $\mathbf{0.82}$ 0.99/0.99/0.99 /$\mathbf{0.79}$ 1.67/1.67/1.67 /$\underline{1.35}$
Durian $23.85$ $44.3$ $\underline{12.48}$ $70.2$ $23.34$ $63.7$ 17.6/17.6/17.6 /$\mathbf{11.43}$ $\underline{42.8}$///$\mathbf{41.1}$
Man $2.52$ $49.0$ $\mathbf{2.18}$ $50.5$ $2.89$ $34.33$ 2.51/2.51/2.51 /$\underline{2.31}$ $\underline{27.2}$///$\mathbf{19.10}$
Sculpt $1.95$ $10.0$ $1.45$ $4.54$ 1.34 $\underline{3.63}$ $\underline{1.14}$///$\mathbf{1.03}$ 4.41/4.41/4.41 /$\mathbf{3.40}$
Stone $10.30$ $43.28$ $\underline{4.35}$ $46.99$ $14.53$ $44.26$ 8.48/8.48/8.48 /$\mathbf{3.93}$ $\underline{27.10}$///$\mathbf{24.73}$
@@
Our experimental results also show that, our method retrieves accurate reconstructed meshes already at lvl $4$, which only contains about $500$K points ($30$ times less than VOXURF). Our method even obtains better results than VOXURF at lvl $4$ for the data Sculpt. Our method is also able to produce significantly fewer artifacts than other methods as shown by the completeness results.
@@ 4.3 Experiments on 4D Human Outfit dataset
We used $8$ scenes of the 4D Human Outfit dataset. This dataset contains sets of $63$ high resolution images with calibrated cameras and ground truth 3D mesh captured with a millimeter precision laser scanner. With this dataset we evaluate the ability of our proposed method to reconstruct detailed geometry such as clothing wrinkles. Note that the mannequins have arms close to the body, which makes it quite favorable to uniform discretization methods like VOXURF. Nevertheless Table 2 shows that our method obtained better results in most scenes. Our proposed method always obtained better results than NeuS2. The advantage of our proposed method is more evident on Figure 6 where we can see significantly better level of details in the face and wrinkles.
@@ 5 Conclusion
We propose a novel method to reconstruct 3D geometry of a target scene from a set of multi-view images by optimizing a SDF field on a Centroidal Voronoi Tesselation. We formulate the optimization framework over the CVT and its dual tetrahedral mesh, designing an efficient framework to output detailed 3D shapes with competitive computation times. Our experimental results validate the key ideas in our proposed method and demonstrate at equivalent discretization level we can achieve a significantly higher level of extracted detail in a majority of situations, compared to competitive approaches. In a number of occurrences our method outperforms or is competitive with SOTA methods while using a magnitude lower level of discretization. Our work opens new promising directions toward detailed 3D reconstruction of large scale scenes under a contained computational time and GPU memory budget.
@@ 6 Acknowledgement
This work was supported by JSPS/KAKENHI JP23H03439 in Japan.

### v2

# VortSDF: 3D Modeling with Centroidal Voronoi Tessellation on Signed Distance Field

@@ Abstract
Volumetric shape representations have become ubiquitous in multi-view reconstruction tasks. They often build on regular voxel grids as discrete representations of 3D shape functions, such as SDF or radiance fields, either as the full shape model or as sampled instantiations of continuous representations, as with neural networks. Despite their proven efficiency, voxel representations come with the precision versus complexity trade-off. This inherent limitation can significantly impact performance when moving away from simple and uncluttered scenes. In this paper we investigate an alternative discretization strategy with the Centroidal Voronoi Tessellation (CVT). CVTs allow to better partition the observation space with respect to shape occupancy and to focus the discretization around shape surfaces. To leverage this discretization strategy for multi-view reconstruction, we introduce a volumetric optimization framework that combines explicit SDF fields with a shallow color network, in order to estimate 3D shape properties over tetrahedral grids. Experimental results with Chamfer statistics validate this approach with unprecedented reconstruction quality on various scenarios such as objects, open scenes or human.
@@ 3.2 Coarse-to-fine Centroidal Voronoi Tesselation
Figure 3: The site locations are optimized using an approximated CVT algorithm that does not explicitly identify the Voronoi cells (left) but consider the neighboring bisectors instead (right).
@@ 3.2.1 SDF-aware implicit CVT optimization
where $b(\mathbf{s_{i}},\mathbf{s_{j}},\mathbf{r})$ is the distance from $\mathbf{s_{i}}$ to its bisector with $\mathbf{s_{j}}$ in the direction $\mathbf{r}$.
@@ 3.4.1 Normal Smoothing
$L_{reg}=0.5\sum_{t}(1-(\frac{\mathbf{\nabla sdf_{t}}\cdot\mathbf{\nabla sdf_{t}}^{smooth}}{\|\mathbf{\nabla sdf_{t}}\|_{2}\|\mathbf{\nabla sdf_{t}}^{smooth}\|_{2}})^{2}).$
@@
where $w_{reg}$ and $w_{tv}$ are weight factors and $L_{TV}$ is a total variation loss.
@@ 4.1 Metrics
We evaluate the quality of the estimated geometry using the available ground truth 3D meshes and a point to mesh Chamfer distance. We compute these errors from the ground truth mesh to the predicted meshes to obtain accuracy $Acc$ [^3] When computing $Acc$, cluttered regions are naturally removed. and from the predicted mesh to the ground truth mesh to obtain completeness $Compl$. We clip the errors to a maximum distance of $0.1$ meter. Note that smaller values are better for these metrics.
@@ 4.2 Experiments on BlendedMVS dataset
We evaluate our method at the last $2$ levels of densification. Our method usually terminates with about $2$M points at lvl $5$ and about $500$K points at lvl 4444444In our experiments we performed up-sampling every $10000$ iterations.. In comparison VOXURF uses a uniform grid of $256\times 256\times 256$ voxels, which corresponds to about $16$M points.
@@
Dog ($31$ images) $1.54$ $19.1$ $1.79$ $\mathbf{4.81}$ $\underline{0.98}$ $11.6$ 1.83/1.83/1.83 /$\mathbf{0.96}$ 18.72/18.72/18.72 /$\underline{9.88}$
Bear ($123$ images) $6.10$ $\mathbf{11.25}$ $\underline{2.07}$ $70.8$ $3.17$ $44.08$ 2.20/2.20/2.20 /$\mathbf{1.41}$ $3/3$$\underline{26.3}$
Clock ($143$ images) $1.34$ $11.8$ $1.11$ $3.29$ $\underline{0.95}$ $\mathbf{0.82}$ 0.99/0.99/0.99 /$\mathbf{0.79}$ 1.67/1.67/1.67 /$\underline{1.35}$
Durian ($124$ images) $23.85$ $44.3$ $\underline{12.48}$ $70.2$ $23.34$ $63.7$ 17.6/17.6/17.6 /$\mathbf{11.43}$ $\underline{42.8}$///$\mathbf{41.1}$
Man ($24$ images) $2.52$ $49.0$ $\mathbf{2.18}$ $50.5$ $2.89$ $34.33$ 2.51/2.51/2.51 /$\underline{2.31}$ $\underline{27.2}$///$\mathbf{19.10}$
Sculpt ($79$ images) $1.95$ $10.0$ $1.45$ $4.54$ 1.34 $\underline{3.63}$ $\underline{1.14}$///$\mathbf{1.03}$ 4.41/4.41/4.41 /$\mathbf{3.40}$
Stone ($56$ images) $10.30$ $43.28$ $\underline{4.35}$ $46.99$ $14.53$ $44.26$ 8.48/8.48/8.48 /$\mathbf{3.93}$ $\underline{27.10}$///$\mathbf{24.73}$
@@
Our experimental results also show that, our method retrieves accurate reconstructed meshes already at lvl $4$, which only contains about $500$K points ($30$ times less than VOXURF). Our method even obtains better results than VOXURF at lvl $4$ for the data Sculpt. Our method is also able to produce significantly fewer artifacts than other methods as shown by the completeness results. Our method converges in about $30$mn at level $4$ and about $50$mn at level $5$. In comparison, NeuS converges in about $8$h30303030mn, NeuS2 in $5$mn and VOXURF in $45$mn.
@@ 4.3 Experiments on 4D Human Outfit dataset
We used $8$ scenes of the 4D Human Outfit dataset. This dataset contains sets of $63$ high resolution images with calibrated cameras and ground truth 3D mesh captured with a millimeter precision laser scanner. With this dataset we evaluate the ability of VortSDF to reconstruct detailed geometry such as clothing wrinkles. Note that the mannequins have arms close to the body, which makes it quite favorable to uniform discretization methods. Yet Table 2 shows that VortSDF obtained better results in most scenes. VortSDF always obtained better results than NeuS2. The advantage VortSDF is more evident on Figure 6 where we can see significantly better level of details in the face and wrinkles.
@@ 5 Conclusion
We propose a novel method to reconstruct 3D geometry of a target scene from a set of multi-view images by optimizing a SDF field on a Centroidal Voronoi Tessellation. We formulate the optimization framework over the CVT and its dual tetrahedral mesh, designing an efficient framework to output detailed 3D shapes with competitive computation times. Our experimental results validate the key ideas in our proposed method and demonstrate at equivalent discretization level we can achieve a significantly higher level of extracted detail in a majority of situations, compared to competitive approaches. In a number of occurrences our method outperforms or is competitive with SOTA methods while using a magnitude lower level of discretization. Our work opens new promising directions toward detailed 3D reconstruction of large scale scenes under a contained computational time and GPU memory budget.
@@ 6 Acknowledgement
This work was in part supported by JSPS/KAKENHI JP23H03439 in Japan.